Find the dimensions of a right triangle where the second leg is 1 unit less than
double the first leg and the hypotenuse is 1 unit more than double the length of
the first leg.

Answer by nyc_function(2741)

second leg = 2x - 1

first leg = x

hypotenuse = 2x + 1

(first leg)^2 + (second leg)^2 = (hypotenuse)^2

x^2 + (2x - 1)^2 = (2x + 1)^2

We get two answers for x: x = 0 and x = 8.

We reject x = 0 because distance must be greater than zero.

The first leg is 8^2 or 64.


The second leg is (2(8) - 1)^2 or 15^2 or 225.

The hypotenuse is (2(8) + 1)^2 = 17^2 or 289.